Find all polar representations, indicate the principle one. 21 = 1 + 6i and z2 = -12i.
Answer #1

Given, z1=1+6i and z2= - 12i = 0 - 12i

First, let us know the general steps to solve this kind of problem. Following that, we will proceed with our problem.See the image below:

polar representation of complex number where, r=121= Ja²b2 a=rcoso bar sino 2-atib b 2=r(coso+isino) and O = tan (ba) and o (

0 = 80.53 180° = r radian 1° = n h 180 8053° x 80.53 nad 180 1.1 rad z, zr(coso tisino) 37 [cos (114) + isen (14) 2

  Z = 2 i 0-12 i a=0, b =-12 144 12 2 [22+62 = (6²+4212 O= tan (42 xo) tanto) = 0o = 0 rad . Z z P(coso tisino) ory 3 = 12 [ c